the segment A(-1,4) to B(2,-2) is extended three times its own length. the terminal point is.
Answer by lwsshak3(11628)

The segment A(-1,4) to B(2,-2) is extended three times its own length. the terminal point is.
***
Extending the segment 3 times its own length means you have a segment divided into 4 equal parts.
Likewise, this means the x-leg and y-leg of the right triangle of which the hypotenuse is the extended segment are also divided into 4 equal parts.
..
length of one x-part=3(-1 to 2)
length of one y-part=6(4 to -2)
..
coordinates of the terminal point:
x= (-1+4*3)=-1+12=11
y=(4+4*6)=4+24=28
(x,y)=(11/28)